OpenStudy (anonymous):

Which of these points lies on the line described by the equation below? y - 8 = 3(x + 6) A). (-8,6) B). (-6,8) C). (6,-8) D). (8,-6)

OpenStudy (anonymous):

i need help

pooja195 (pooja195):

Same thing as the previous ones y - 8 = 3(x + 6) y-8=3x+18 y=3x+26 Now plug in the given points and see which statement is true 6=3(-8)+26 8=3(-6)+26 -8=3(6)+26 -6=3(8)+26

OpenStudy (anonymous):

i need answers